# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ Appendix E Full Results of Our Proposed Practical Setting
Strategy Channel-Dependent Channel-Independent Models CTF(ours) TimeXer TimeMixer iTrans. CrossGNN TimesNet PatchTST DLinear Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE $m=0.125$ 288 0.331 $\pm$ 0.004 0.412 $\pm$ 0.003 0.355 $\pm$ 0.058 0.438 $\pm$ 0.050 0.372 $\pm$ 0.012 0.456 $\pm$ 0.023 0.368 $\pm$ 0.005 0.432 $\pm$ 0.005 0.358 $\pm$ 0.013 0.441 $\pm$ 0.012 0.353 $\pm$ 0.009 0.416 $\pm$ 0.006 0.372 $\pm$ 0.017 0.457 $\pm$ 0.010 0.352 $\pm$ 0.004 0.464 $\pm$ 0.004 576 0.393 $\pm$ 0.002 0.456 $\pm$ 0.001 0.434 $\pm$ 0.042 0.498 $\pm$ 0.038 0.452 $\pm$ 0.013 0.508 $\pm$ 0.013 0.451 $\pm$ 0.003 0.487 $\pm$ 0.003 0.445 $\pm$ 0.010 0.496 $\pm$ 0.008 0.494 $\pm$ 0.038 0.497 $\pm$ 0.026 0.454 $\pm$ 0.017 0.522 $\pm$ 0.012 0.428 $\pm$ 0.002 0.532 $\pm$ 0.003 864 0.418 $\pm$ 0.008 0.475 $\pm$ 0.007 0.444 $\pm$ 0.061 0.495 $\pm$ 0.017 0.474 $\pm$ 0.016 0.514 $\pm$ 0.022 0.481 $\pm$ 0.015 0.506 $\pm$ 0.006 0.505 $\pm$ 0.009 0.533 $\pm$ 0.005 0.474 $\pm$ 0.026 0.496 $\pm$ 0.016 0.486 $\pm$ 0.022 0.544 $\pm$ 0.016 0.466 $\pm$ 0.004 0.563 $\pm$ 0.004 1152 0.441 $\pm$ 0.005 0.489 $\pm$ 0.005 0.495 $\pm$ 0.032 0.533 $\pm$ 0.027 0.480 $\pm$ 0.027 0.519 $\pm$ 0.022 0.538 $\pm$ 0.006 0.534 $\pm$ 0.003 0.560 $\pm$ 0.011 0.566 $\pm$ 0.014 0.537 $\pm$ 0.011 0.529 $\pm$ 0.011 0.499 $\pm$ 0.019 0.538 $\pm$ 0.015 0.480 $\pm$ 0.001 0.575 $\pm$ 0.001 $m=0.25$ 288 0.347 $\pm$ 0.005 0.433 $\pm$ 0.003 0.386 $\pm$ 0.059 0.465 $\pm$ 0.048 0.389 $\pm$ 0.012 0.482 $\pm$ 0.023 0.387 $\pm$ 0.004 0.455 $\pm$ 0.004 0.375 $\pm$ 0.014 0.468 $\pm$ 0.013 0.371 $\pm$ 0.009 0.445 $\pm$ 0.006 0.430 $\pm$ 0.022 0.514 $\pm$ 0.011 0.374 $\pm$ 0.003 0.489 $\pm$ 0.004 576 0.404 $\pm$ 0.003 0.473 $\pm$ 0.001 0.469 $\pm$ 0.044 0.526 $\pm$ 0.036 0.472 $\pm$ 0.014 0.536 $\pm$ 0.013 0.472 $\pm$ 0.005 0.512 $\pm$ 0.004 0.455 $\pm$ 0.010 0.523 $\pm$ 0.008 0.523 $\pm$ 0.041 0.528 $\pm$ 0.025 0.533 $\pm$ 0.025 0.593 $\pm$ 0.018 0.447 $\pm$ 0.002 0.555 $\pm$ 0.002 864 0.434 $\pm$ 0.006 0.495 $\pm$ 0.005 0.496 $\pm$ 0.043 0.530 $\pm$ 0.012 0.484 $\pm$ 0.020 0.534 $\pm$ 0.026 0.504 $\pm$ 0.013 0.528 $\pm$ 0.008 0.509 $\pm$ 0.009 0.559 $\pm$ 0.004 0.481 $\pm$ 0.031 0.517 $\pm$ 0.017 0.573 $\pm$ 0.037 0.625 $\pm$ 0.021 0.482 $\pm$ 0.003 0.583 $\pm$ 0.003 1152 0.455 $\pm$ 0.008 0.508 $\pm$ 0.005 0.542 $\pm$ 0.021 0.564 $\pm$ 0.018 0.510 $\pm$ 0.023 0.536 $\pm$ 0.029 0.555 $\pm$ 0.006 0.557 $\pm$ 0.003 0.557 $\pm$ 0.013 0.590 $\pm$ 0.016 0.554 $\pm$ 0.016 0.554 $\pm$ 0.014 0.571 $\pm$ 0.043 0.607 $\pm$ 0.029 0.496 $\pm$ 0.001 0.596 $\pm$ 0.001 $m=0.375$ 288 0.363 $\pm$ 0.004 0.458 $\pm$ 0.002 0.412 $\pm$ 0.058 0.487 $\pm$ 0.048 0.421 $\pm$ 0.013 0.510 $\pm$ 0.023 0.415 $\pm$ 0.004 0.483 $\pm$ 0.005 0.400 $\pm$ 0.016 0.498 $\pm$ 0.014 0.398 $\pm$ 0.009 0.482 $\pm$ 0.008 0.492 $\pm$ 0.022 0.551 $\pm$ 0.016 0.400 $\pm$ 0.003 0.515 $\pm$ 0.003 576 0.418 $\pm$ 0.004 0.494 $\pm$ 0.003 0.498 $\pm$ 0.036 0.548 $\pm$ 0.025 0.509 $\pm$ 0.020 0.566 $\pm$ 0.015 0.501 $\pm$ 0.007 0.539 $\pm$ 0.004 0.473 $\pm$ 0.009 0.552 $\pm$ 0.007 0.538 $\pm$ 0.036 0.557 $\pm$ 0.020 0.612 $\pm$ 0.034 0.650 $\pm$ 0.020 0.470 $\pm$ 0.002 0.579 $\pm$ 0.002 864 0.445 $\pm$ 0.011 0.516 $\pm$ 0.006 0.542 $\pm$ 0.021 0.553 $\pm$ 0.008 0.511 $\pm$ 0.024 0.560 $\pm$ 0.026 0.536 $\pm$ 0.010 0.556 $\pm$ 0.008 0.521 $\pm$ 0.003 0.585 $\pm$ 0.004 0.499 $\pm$ 0.031 0.547 $\pm$ 0.016 0.651 $\pm$ 0.046 0.685 $\pm$ 0.030 0.501 $\pm$ 0.003 0.605 $\pm$ 0.002 1152 0.472 $\pm$ 0.008 0.532 $\pm$ 0.004 0.579 $\pm$ 0.013 0.587 $\pm$ 0.016 0.556 $\pm$ 0.023 0.565 $\pm$ 0.030 0.574 $\pm$ 0.009 0.584 $\pm$ 0.006 0.564 $\pm$ 0.021 0.615 $\pm$ 0.022 0.575 $\pm$ 0.028 0.585 $\pm$ 0.020 0.644 $\pm$ 0.061 0.667 $\pm$ 0.042 0.516 $\pm$ 0.001 0.618 $\pm$ 0.001 $m=0.5$ 288 0.390 $\pm$ 0.004 0.488 $\pm$ 0.003 0.445 $\pm$ 0.052 0.508 $\pm$ 0.056 0.456 $\pm$ 0.015 0.533 $\pm$ 0.022 0.460 $\pm$ 0.007 0.508 $\pm$ 0.005 0.429 $\pm$ 0.014 0.526 $\pm$ 0.011 0.421 $\pm$ 0.010 0.514 $\pm$ 0.008 0.566 $\pm$ 0.026 0.571 $\pm$ 0.012 0.428 $\pm$ 0.002 0.540 $\pm$ 0.003 576 0.444 $\pm$ 0.003 0.525 $\pm$ 0.002 0.528 $\pm$ 0.028 0.568 $\pm$ 0.015 0.542 $\pm$ 0.025 0.585 $\pm$ 0.014 0.550 $\pm$ 0.011 0.568 $\pm$ 0.004 0.496 $\pm$ 0.007 0.581 $\pm$ 0.005 0.574 $\pm$ 0.036 0.592 $\pm$ 0.018 0.674 $\pm$ 0.034 0.685 $\pm$ 0.022 0.494 $\pm$ 0.001 0.602 $\pm$ 0.002 864 0.470 $\pm$ 0.010 0.544 $\pm$ 0.005 0.586 $\pm$ 0.006 0.574 $\pm$ 0.013 0.548 $\pm$ 0.023 0.587 $\pm$ 0.023 0.597 $\pm$ 0.011 0.589 $\pm$ 0.006 0.538 $\pm$ 0.006 0.613 $\pm$ 0.007 0.528 $\pm$ 0.029 0.580 $\pm$ 0.014 0.708 $\pm$ 0.036 0.713 $\pm$ 0.023 0.522 $\pm$ 0.002 0.626 $\pm$ 0.001 1152 0.506 $\pm$ 0.006 0.563 $\pm$ 0.002 0.612 $\pm$ 0.018 0.607 $\pm$ 0.016 0.612 $\pm$ 0.024 0.594 $\pm$ 0.028 0.617 $\pm$ 0.011 0.615 $\pm$ 0.005 0.577 $\pm$ 0.035 0.642 $\pm$ 0.028 0.611 $\pm$ 0.038 0.621 $\pm$ 0.023 0.704 $\pm$ 0.071 0.703 $\pm$ 0.037 0.538 $\pm$ 0.001 0.640 $\pm$ 0.001
@@
Train missing ratio
@@
0.548
0.476
0.540
0.452
0.530
0.456
0.528
Table 16: Full results of the ETT1 dataset under short-range test-time missing intervals with varying missing ratios $m$. CTF is trained with a random patch masking ratio of 0.4 to enhance robustness against missing inputs.
Strategy

### v2

@@ Appendix E Full Results of Our Proposed Practical Setting
Approach Channel-Dependent Channel-Independent Irregular modeling Missing Model CTF(ours) TimeXer iTrans. CrossGNN TimesNet TimeMixer++ PatchTST DLinear Hi-Patch tPatchGNN ContiFormer BiTGraph Metric CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE CMSE CMAE $m=0.125$ 288 0.336 $\pm$ 0.003 0.414 $\pm$ 0.003 0.344 $\pm$ 0.010 0.419 $\pm$ 0.011 0.367 $\pm$ 0.006 0.426 $\pm$ 0.004 0.347 $\pm$ 0.011 0.412 $\pm$ 0.015 0.366 $\pm$ 0.011 0.409 $\pm$ 0.008 0.360 $\pm$ 0.013 0.415 $\pm$ 0.007 0.337 $\pm$ 0.007 0.413 $\pm$ 0.005 0.349 $\pm$ 0.005 0.458 $\pm$ 0.004 0.369 $\pm$ 0.007 0.441 $\pm$ 0.008 0.380 $\pm$ 0.014 0.454 $\pm$ 0.011 0.353 $\pm$ 0.002 0.426 $\pm$ 0.004 0.333 $\pm$ 0.002 0.429 $\pm$ 0.004 576 0.400 $\pm$ 0.002 0.456 $\pm$ 0.002 0.427 $\pm$ 0.006 0.477 $\pm$ 0.008 0.459 $\pm$ 0.004 0.484 $\pm$ 0.004 0.445 $\pm$ 0.010 0.472 $\pm$ 0.010 0.478 $\pm$ 0.027 0.479 $\pm$ 0.018 0.433 $\pm$ 0.009 0.470 $\pm$ 0.008 0.423 $\pm$ 0.011 0.479 $\pm$ 0.007 0.423 $\pm$ 0.003 0.522 $\pm$ 0.003 0.429 $\pm$ 0.023 0.475 $\pm$ 0.021 0.448 $\pm$ 0.015 0.501 $\pm$ 0.017 0.429 $\pm$ 0.006 0.477 $\pm$ 0.004 0.413 $\pm$ 0.028 0.497 $\pm$ 0.028 864 0.433 $\pm$ 0.005 0.483 $\pm$ 0.005 0.457 $\pm$ 0.020 0.492 $\pm$ 0.016 0.517 $\pm$ 0.018 0.518 $\pm$ 0.007 0.515 $\pm$ 0.009 0.514 $\pm$ 0.007 0.507 $\pm$ 0.014 0.502 $\pm$ 0.010 0.466 $\pm$ 0.016 0.498 $\pm$ 0.011 0.462 $\pm$ 0.019 0.508 $\pm$ 0.016 0.457 $\pm$ 0.005 0.549 $\pm$ 0.004 0.443 $\pm$ 0.013 0.484 $\pm$ 0.011 0.471 $\pm$ 0.021 0.512 $\pm$ 0.023 0.472 $\pm$ 0.003 0.513 $\pm$ 0.002 0.447 $\pm$ 0.031 0.526 $\pm$ 0.034 1152 0.465 $\pm$ 0.002 0.500 $\pm$ 0.003 0.482 $\pm$ 0.013 0.508 $\pm$ 0.011 0.559 $\pm$ 0.007 0.538 $\pm$ 0.004 0.581 $\pm$ 0.010 0.553 $\pm$ 0.011 0.523 $\pm$ 0.062 0.507 $\pm$ 0.039 0.484 $\pm$ 0.016 0.510 $\pm$ 0.009 0.481 $\pm$ 0.007 0.507 $\pm$ 0.004 0.475 $\pm$ 0.001 0.565 $\pm$ 0.001 0.466 $\pm$ 0.009 0.494 $\pm$ 0.004 0.502 $\pm$ 0.008 0.520 $\pm$ 0.006 0.512 $\pm$ 0.003 0.535 $\pm$ 0.002 0.510 $\pm$ 0.062 0.585 $\pm$ 0.058 $m=0.25$ 288 0.350 $\pm$ 0.001 0.431 $\pm$ 0.002 0.358 $\pm$ 0.011 0.434 $\pm$ 0.012 0.389 $\pm$ 0.007 0.447 $\pm$ 0.006 0.359 $\pm$ 0.010 0.428 $\pm$ 0.012 0.379 $\pm$ 0.012 0.423 $\pm$ 0.004 0.389 $\pm$ 0.016 0.438 $\pm$ 0.007 0.361 $\pm$ 0.010 0.439 $\pm$ 0.006 0.361 $\pm$ 0.004 0.473 $\pm$ 0.005 0.381 $\pm$ 0.007 0.460 $\pm$ 0.010 0.390 $\pm$ 0.015 0.467 $\pm$ 0.012 0.368 $\pm$ 0.003 0.451 $\pm$ 0.005 0.343 $\pm$ 0.008 0.438 $\pm$ 0.007 576 0.417 $\pm$ 0.002 0.473 $\pm$ 0.001 0.440 $\pm$ 0.006 0.490 $\pm$ 0.008 0.479 $\pm$ 0.005 0.505 $\pm$ 0.004 0.459 $\pm$ 0.009 0.489 $\pm$ 0.008 0.511 $\pm$ 0.022 0.505 $\pm$ 0.016 0.463 $\pm$ 0.011 0.492 $\pm$ 0.009 0.454 $\pm$ 0.016 0.514 $\pm$ 0.010 0.434 $\pm$ 0.003 0.536 $\pm$ 0.002 0.459 $\pm$ 0.031 0.499 $\pm$ 0.022 0.463 $\pm$ 0.019 0.518 $\pm$ 0.020 0.441 $\pm$ 0.005 0.504 $\pm$ 0.004 0.428 $\pm$ 0.035 0.511 $\pm$ 0.038 864 0.456 $\pm$ 0.006 0.503 $\pm$ 0.005 0.472 $\pm$ 0.018 0.507 $\pm$ 0.015 0.537 $\pm$ 0.018 0.537 $\pm$ 0.010 0.531 $\pm$ 0.007 0.531 $\pm$ 0.005 0.487 $\pm$ 0.045 0.487 $\pm$ 0.020 0.500 $\pm$ 0.020 0.525 $\pm$ 0.012 0.496 $\pm$ 0.014 0.546 $\pm$ 0.012 0.467 $\pm$ 0.005 0.564 $\pm$ 0.004 0.467 $\pm$ 0.012 0.505 $\pm$ 0.012 0.492 $\pm$ 0.024 0.531 $\pm$ 0.030 0.487 $\pm$ 0.002 0.544 $\pm$ 0.003 0.465 $\pm$ 0.028 0.542 $\pm$ 0.034 1152 0.491 $\pm$ 0.006 0.521 $\pm$ 0.003 0.498 $\pm$ 0.011 0.521 $\pm$ 0.011 0.580 $\pm$ 0.007 0.558 $\pm$ 0.004 0.599 $\pm$ 0.009 0.572 $\pm$ 0.013 0.562 $\pm$ 0.029 0.533 $\pm$ 0.015 0.518 $\pm$ 0.019 0.538 $\pm$ 0.016 0.511 $\pm$ 0.010 0.538 $\pm$ 0.007 0.484 $\pm$ 0.002 0.578 $\pm$ 0.001 0.496 $\pm$ 0.010 0.516 $\pm$ 0.008 0.523 $\pm$ 0.009 0.540 $\pm$ 0.007 0.529 $\pm$ 0.003 0.566 $\pm$ 0.003 0.537 $\pm$ 0.056 0.602 $\pm$ 0.055 $m=0.375$ 288 0.370 $\pm$ 0.004 0.457 $\pm$ 0.002 0.373 $\pm$ 0.016 0.450 $\pm$ 0.019 0.436 $\pm$ 0.008 0.487 $\pm$ 0.006 0.381 $\pm$ 0.008 0.451 $\pm$ 0.009 0.442 $\pm$ 0.071 0.469 $\pm$ 0.041 0.451 $\pm$ 0.018 0.482 $\pm$ 0.007 0.406 $\pm$ 0.015 0.479 $\pm$ 0.009 0.378 $\pm$ 0.004 0.495 $\pm$ 0.004 0.384 $\pm$ 0.010 0.473 $\pm$ 0.009 0.410 $\pm$ 0.020 0.491 $\pm$ 0.014 0.392 $\pm$ 0.005 0.477 $\pm$ 0.006 0.362 $\pm$ 0.016 0.453 $\pm$ 0.012 576 0.439 $\pm$ 0.003 0.499 $\pm$ 0.002 0.457 $\pm$ 0.015 0.505 $\pm$ 0.016 0.521 $\pm$ 0.006 0.541 $\pm$ 0.006 0.483 $\pm$ 0.006 0.512 $\pm$ 0.006 0.533 $\pm$ 0.019 0.527 $\pm$ 0.012 0.522 $\pm$ 0.010 0.536 $\pm$ 0.007 0.518 $\pm$ 0.021 0.568 $\pm$ 0.014 0.449 $\pm$ 0.003 0.557 $\pm$ 0.003 0.470 $\pm$ 0.022 0.514 $\pm$ 0.024 0.491 $\pm$ 0.031 0.544 $\pm$ 0.027 0.458 $\pm$ 0.005 0.531 $\pm$ 0.006 0.448 $\pm$ 0.038 0.527 $\pm$ 0.039 864 0.481 $\pm$ 0.007 0.528 $\pm$ 0.005 0.496 $\pm$ 0.016 0.524 $\pm$ 0.011 0.580 $\pm$ 0.016 0.570 $\pm$ 0.010 0.556 $\pm$ 0.005 0.555 $\pm$ 0.003 0.562 $\pm$ 0.016 0.545 $\pm$ 0.010 0.562 $\pm$ 0.022 0.572 $\pm$ 0.016 0.565 $\pm$ 0.017 0.604 $\pm$ 0.011 0.482 $\pm$ 0.006 0.584 $\pm$ 0.004 0.494 $\pm$ 0.014 0.527 $\pm$ 0.013 0.530 $\pm$ 0.022 0.561 $\pm$ 0.027 0.505 $\pm$ 0.004 0.574 $\pm$ 0.004 0.488 $\pm$ 0.027 0.561 $\pm$ 0.028 1152 0.516 $\pm$ 0.008 0.547 $\pm$ 0.004 0.520 $\pm$ 0.012 0.540 $\pm$ 0.010 0.620 $\pm$ 0.009 0.592 $\pm$ 0.005 0.626 $\pm$ 0.009 0.596 $\pm$ 0.013 0.611 $\pm$ 0.016 0.568 $\pm$ 0.015 0.588 $\pm$ 0.032 0.591 $\pm$ 0.020 0.573 $\pm$ 0.017 0.591 $\pm$ 0.011 0.498 $\pm$ 0.001 0.598 $\pm$ 0.001 0.524 $\pm$ 0.014 0.538 $\pm$ 0.010 0.553 $\pm$ 0.014 0.570 $\pm$ 0.012 0.546 $\pm$ 0.005 0.600 $\pm$ 0.005 0.570 $\pm$ 0.058 0.626 $\pm$ 0.053 $m=0.5$ 288 0.393 $\pm$ 0.002 0.483 $\pm$ 0.001 0.419 $\pm$ 0.011 0.488 $\pm$ 0.010 0.511 $\pm$ 0.013 0.542 $\pm$ 0.009 0.418 $\pm$ 0.006 0.484 $\pm$ 0.005 0.477 $\pm$ 0.022 0.498 $\pm$ 0.008 0.561 $\pm$ 0.032 0.550 $\pm$ 0.010 0.469 $\pm$ 0.022 0.526 $\pm$ 0.012 0.407 $\pm$ 0.004 0.525 $\pm$ 0.005 0.405 $\pm$ 0.010 0.498 $\pm$ 0.010 0.430 $\pm$ 0.016 0.516 $\pm$ 0.011 0.423 $\pm$ 0.009 0.500 $\pm$ 0.005 0.387 $\pm$ 0.024 0.470 $\pm$ 0.018 576 0.462 $\pm$ 0.003 0.523 $\pm$ 0.002 0.510 $\pm$ 0.004 0.542 $\pm$ 0.007 0.587 $\pm$ 0.008 0.592 $\pm$ 0.007 0.523 $\pm$ 0.005 0.545 $\pm$ 0.002 0.607 $\pm$ 0.041 0.568 $\pm$ 0.030 0.623 $\pm$ 0.023 0.602 $\pm$ 0.012 0.599 $\pm$ 0.027 0.624 $\pm$ 0.017 0.477 $\pm$ 0.003 0.586 $\pm$ 0.003 0.498 $\pm$ 0.020 0.542 $\pm$ 0.023 0.513 $\pm$ 0.044 0.571 $\pm$ 0.036 0.487 $\pm$ 0.006 0.557 $\pm$ 0.007 0.473 $\pm$ 0.038 0.548 $\pm$ 0.054 864 0.505 $\pm$ 0.007 0.554 $\pm$ 0.005 0.550 $\pm$ 0.016 0.563 $\pm$ 0.015 0.645 $\pm$ 0.021 0.620 $\pm$ 0.012 0.599 $\pm$ 0.005 0.589 $\pm$ 0.005 0.621 $\pm$ 0.016 0.580 $\pm$ 0.013 0.662 $\pm$ 0.027 0.642 $\pm$ 0.021 0.651 $\pm$ 0.026 0.665 $\pm$ 0.014 0.508 $\pm$ 0.007 0.613 $\pm$ 0.004 0.533 $\pm$ 0.027 0.558 $\pm$ 0.017 0.563 $\pm$ 0.014 0.596 $\pm$ 0.019 0.539 $\pm$ 0.007 0.603 $\pm$ 0.006 0.515 $\pm$ 0.024 0.583 $\pm$ 0.028 1152 0.538 $\pm$ 0.007 0.572 $\pm$ 0.004 0.577 $\pm$ 0.012 0.582 $\pm$ 0.015 0.681 $\pm$ 0.010 0.638 $\pm$ 0.007 0.673 $\pm$ 0.016 0.631 $\pm$ 0.020 0.675 $\pm$ 0.030 0.609 $\pm$ 0.015 0.688 $\pm$ 0.055 0.661 $\pm$ 0.029 0.653 $\pm$ 0.024 0.648 $\pm$ 0.016 0.522 $\pm$ 0.001 0.625 $\pm$ 0.001 0.565 $\pm$ 0.022 0.570 $\pm$ 0.017 0.583 $\pm$ 0.016 0.602 $\pm$ 0.015 0.582 $\pm$ 0.009 0.631 $\pm$ 0.007 0.606 $\pm$ 0.068 0.649 $\pm$ 0.049
@@
Missing ratio (Train)
@@
0.541
0.474
0.533
0.474
0.536
Table 19: Average CMSE and CMAE for all prediction lengths of ETT1 dataset under short-range test-time missing intervals with varying missing ratios $m$. CTF is trained with a random patch masking ratio of 0.4 to enhance robustness against missing inputs.
Approach
@@
Table 20: Average CMSE and CMAE for all prediction lengths of EPA dataset (LA, Hillsborough) under short-range test-time missing intervals with varying missing ratios $m$. CTF is trained with a random patch masking ratio of 0.4 to enhance robustness against missing inputs.
Model CTF (ours) Hi-Patch ContiFormer
Metric CMSE CMAE CMSE CMAE CMSE CMAE
$m=0.1$ 1.074 0.702 1.084 0.742 1.219 0.744
$m=0.2$ 1.146 0.731 1.121 0.754 1.237 0.757
$m=0.3$ 1.226 0.756 1.204 0.781 1.254 0.773
$m=0.4$ 1.290 0.788 1.383 0.839 1.298 0.801